### v1

@@ 4.2 Estimation strategy
where $\mathcal{D}$ is the Kullback-Leibler divergence customized to the driven autoregressive Poisson model as described in Equations (15) and (19), $\textbf{D}_{2}$ is the discrete Laplacian, which, combined with the $\ell_{1}$ norms favors piecewise linear behavior of the estimate, and $\lambda>0$ is a regularization parameter balancing the data-fidelity and the penalization [1, 49]. As the minimization problem (43) is convex but nonsmooth, it is solved using the Chambolle-Pock proximal algorithm [9, 1, 49] leveraging the closed-form expression of the proximity operator of the Kullback-Leibler divergence of Equation (15[1].
@@
Let $\boldsymbol{\mathsf{Y}}$ observations following the driven autoregressive Poisson model described in Section 4.1. Then, the variational estimator $\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda)$ defined in Equation (43) satisfies the assumptions enunciated in Theorems 1 and 2. Hence, for any $N\in\mathbb{N}^{*}$, the data-driven robustified oracles $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}$ and $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}$ are asymptotically unbiased estimators of the estimation and prediction risks respectively.
@@
$\begin{split}\mathcal{E}^{\circ}(\boldsymbol{\mathsf{Y}};\lambda)&=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda)-\overline{\boldsymbol{\mathsf{X}}}\rVert^{2}_{2},\\ \mathcal{P}^{\circ}(\boldsymbol{\mathsf{Y}};\lambda)&=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda)\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})-\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\rVert^{2}_{2}\end{split}$ (44)
@@
Metric Minimized oracle $\log_{10}\alpha=2$ $\log_{10}\alpha=2.5$ $\log_{10}\alpha=3$ $\log_{10}\alpha=3.5$ $\log_{10}\alpha=4$ $\log_{10}\alpha=4.5$ $\log_{10}\alpha=5$
$\mathsf{MMSE}\pm\mathsf{CI}$ Estimation error $\mathcal{E}^{\circ}$ $0.51\pm 0.03$ $0.59\pm 0.07$ $1.14\pm 0.36$ $1.53\pm 0.76$ $1.69\pm 0.73$ $1.28\pm 0.36$ $4.20\pm 3.35$
Prediction error $\mathcal{P}^{\circ}$ $0.52\pm 0.03$ $0.61\pm 0.07$ $1.35\pm 0.62$ $1.71\pm 0.74$ $1.85\pm 0.76$ $1.39\pm 0.38$ $5.22\pm 4.01$
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ $0.52\pm 0.03$ $0.62\pm 0.07$ $1.37\pm 0.63$ $1.82\pm 0.74$ $1.84\pm 0.80$ $1.46\pm 0.43$ $20.03\pm 26.27$
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ $0.52\pm 0.03$ $0.61\pm 0.07$ $1.57\pm 0.77$ $6.39\pm 7.31$ $6.98\pm 6.25$ $16.69\pm 19.51$ $209.61\pm 286.66$
$\mathsf{Bias}$$\mathsf{Variance}$ Estimation error $\mathcal{E}^{\circ}$ 0.490.490.490.490.020.020.020.02 0.520.520.520.520.070.070.070.07 0.760.760.760.760.380.380.380.38 0.780.780.780.780.750.750.750.75 0.850.850.850.850.840.840.840.84 0.530.530.530.530.750.750.750.75 1.591.591.591.592.612.612.612.61
Prediction error $\mathcal{P}^{\circ}$ 0.490.490.490.490.030.030.030.03 0.510.510.510.510.090.090.090.09 0.640.640.640.640.710.710.710.71 0.720.720.720.720.990.990.990.99 0.760.760.760.761.081.081.081.08 0.480.480.480.480.910.910.910.91 1.111.111.111.114.114.114.114.11
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.500.500.020.020.020.02 0.540.540.540.540.070.070.070.07 0.680.680.680.680.690.690.690.69 0.750.750.750.751.071.071.071.07 0.970.970.970.970.870.870.870.87 0.590.590.590.590.870.870.870.87 2.672.672.672.6717.3617.3617.3617.36
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.500.500.030.030.030.03 0.580.580.580.580.080.080.080.08 0.660.660.660.660.910.910.910.91 1.161.161.161.165.235.235.235.23 1.281.281.281.285.695.695.695.69 1.911.911.911.9114.7814.7814.7814.78 25.8125.8125.8125.81183.78183.78183.78183.78
@@ 4.3 Performance evaluation
For each scale parameter explored, performances are evaluated on a collection of $Q=10$ independent synthetic observations $\{\boldsymbol{\mathsf{Y}}^{(q)},\,q=1,\ldots,Q\}$ generated following the setup of Section 4.1. For $\mathcal{O}\in\{\mathcal{E}^{\circ},\mathcal{P}^{\circ},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\}$, the hyperparameters selected through the oracle strategy applied to the $q$th realization are denoted [^6] The dependency of $\lambda_{\mathcal{O}}$ in $\boldsymbol{\mathsf{Y}}$ is omitted for readability.
@@
Several performance criteria are considered. First, the Minimal Mean Squared Error over the $Q$ realizations, defined as
@@
$\mathsf{Bias}$ $=\left\lVert\langle\widehat{\boldsymbol{\mathsf{X}}}\rangle_{\mathcal{O}}-\overline{\boldsymbol{\mathsf{X}}}\right\rVert^{2}_{2},$ (48)
$\mathsf{Variance}$ $=\frac{1}{Q}\sum_{q=1}^{Q}\left\lVert\widehat{\boldsymbol{\mathsf{X}}}\left(\boldsymbol{\mathsf{Y}}^{(q)};\lambda_{\mathcal{O}}^{(q)}\right)-\langle\widehat{\boldsymbol{\mathsf{X}}}\rangle_{\mathcal{O}}\right\rVert^{2}$ (49)
@@ 4.4 Results
Figure 2 compares the four oracle-based estimation strategies on one realization of driven autoregressive Poisson synthetic observations for three values of the scale parameter, corresponding to low, medium and high noise levels. Synthetic observations and memory terms are displayed in the first row; the underlying ground truth reproduction coefficient and its estimates obtained using the four different oracles are plotted in blue on the second row; third (resp. fourth) row compares the prediction (resp. estimation) ground truth and data-dependent oracles, and the associated optimal hyperparameters.
@@ 5.1 Weekly scaled Poisson epidemiological model
The considered epidemiological model, briefly introduction in Section 1, in Equation (5), was initially proposed in [15] and states that, conditionally to past infection counts $\mathsf{Z}_{1},\ldots,\mathsf{Z}_{t-1}$, the number of new infections at time $t$, denoted $\mathsf{Z}_{t}$, follows a Poisson distribution of intensity equal to the product of the effective reproduction number at time $t$, $\mathsf{R}_{t}$, and the global infectiousness $\Phi_{t}(\boldsymbol{\mathsf{Z}})=\sum_{s\geq 1}\varphi_{s}\mathsf{Z}_{t-s}$ with $\varphi$ the serial interval distribution. [^8] The serial interval is the random delay between primary and secondary infections; its distribution encodes the typical time scales of the propagation. Although very accurate for a posteriori consolidated data [15], the daily version of this model appeared not suitable for real-time COVID-19 daily infection counts, an example of which is displayed in Figure 3, gray curve. Indeed, COVID-19 data are severely corrupted by administrative noise, taking the form of missing counts during week-ends and holidays, large cumulative counts on Mondays, pseudo-seasonalities, erroneous samples, to name but a few. This noise has two major effects: first, reported COVID-19 daily infections counts contain outlier values, notably during week-ends as illustrated in the gray curve in Figure 3; second, the variance of the reported counts reflects both the intrinsic variance of the propagation process and the additional variance induced by the fluctuating administrative delays and errors. To circumvent the presence of outlier samples, a classical strategy in epidemiology is to consider aggregated data [21, 11, 44], e.g., at the scale of the week as illustrated in Figure 3, black curve, which is far smoother and hence more realistic from an epidemiological point of view. The present work proposes to account for the increased variance observed in real COVID-19 infection counts through a constant in time scale parameter $\alpha>1$. Up to the authors knowledge, scaled Poisson distributions have not yet been explored in the epidemiology literature, making this an original contribution of the present work. Altogether, the proposed weekly scaled Poisson epidemiological model writes
@@
where time instants $t$ corresponds to weeks and the serial interval distribution used to compute $\Phi_{t}(\boldsymbol{\mathsf{Z}})$ is weekly discretized. The weekly discretized COVID-19 serial interval distribution can be obtained by coarsening the daily discretized distribution provided in [8, 55] into a weekly distribution, e.g., by performing an integration over one-week windows using the left rectangle method with a one-day integration step. Under Model (50), the expected number of infections at week $t$ is $\Phi_{t}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{t}$, unchanged compared to the standard Poisson model of Equation (5), but the variance is $\alpha\Phi_{t}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{t}$, larger by a factor $\alpha$ than the variance of the standard Poisson model.
Figure 3: COVID-19 daily vs. weekly new infections counts. French counts from Johns Hopkins University repository, covering 71 weeks from October 4, 2021 to February 6, 2023.
(a) France
(b) India
(c) Canada
(d) Argentina
Figure 4: Data-driven estimation of the reproduction number of COVID-19 from weekly infection counts. Top row: weekly aggregated infection counts. Middle row: estimated effective reproduction number with data-driven hyperparameter selection based on the robustified $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$. Bottom row: robustified autoregressive unbiased prediction risk estimate averaged over $N=10$ realizations of the Monte Carlo vector with associated $95\%$ Gaussian confidence interval for a logarithmically spaced range of regularization parameters and minimizing regularization parameter obtained through grid search.
@@ 5.2 Data-driven reproduction number estimation strategy
leveraging the robustified unbiased prediction risk estimate, whose accuracy and robustness have been demonstrated on synthetic data in Section 4.
@@ 6 Conclusion and perspectives
A novel driven autoregressive model has been proposed, inspired by state-of-the-art viral epidemics models. The estimation of the parameters of the model, namely the time-varying reproduction coefficient, has been framed as a nonstandard, highly nonstationary, inverse problem. The proposed rigorous mathematical formulation enabled to leverage the efficient and versatile variational framework to design accurate reproduction coefficient estimators which are robust to high noise levels in the observations. The third major contribution consisted in the design of asymptotically unbiased risk estimates, which are then plugged into an oracle strategy for fully data-driven fine-tuning of the hyperparameters of the variational estimator, removing the major obstacle to its practical use. The resulting data-driven estimation strategy has been assessed through intensive Monte Carlo simulations on synthetic data. Taking advantage of the proposed extended mathematical framework, the standard epidemiological model for viral epidemics has been enriched to account for the low quality of COVID-19 data, leading to a novel weekly scaled Poisson model. Finally, the data-driven estimation procedure is shown to yield very consistent estimation of the COVID-19 reproduction number in various countries and pandemic stages despite the low quality of reported data, demonstrating its practical applicability for epidemic monitoring an a crisis context. The data-driven nature of the proposed epidemiological indicator estimation strategy constitutes a major asset for its dissemination beyond signal processing as its use requires no expert knowledge.
Further work will consists in, first, extending the unbiased risk estimates to other noise models, e.g., to additive Gaussian and multiplicative Gamma noises inspiring from the generalized Stein estimators [20], in order to further enrich the proposed framework and to be able to extend the developed methodology to other models, in epidemiology or beyond. Second, as in epidemiology the memory functions are often parametric, with parameters encapsulating the pathogen transmission characteristics, the proposed framework will be leveraged to perform simultaneously the extraction of the memory functions parameters and the fine-tuning of the estimator hyperparameters.

### v3

@@ 4.2 Estimation strategy
where $\mathcal{D}$ is the Kullback-Leibler divergence customized to the generalized time-varying autoregressive Poisson model as described in Equation (18), $\textbf{D}_{2}$ is the discrete Laplacian, which, combined with the $\ell_{1}$ norms favors piecewise linear behavior of the estimate, and $\lambda>0$ is a regularization parameter balancing the data-fidelity and the penalization [1, 59]. As the minimization problem (41) is convex but nonsmooth, it is solved using the Chambolle-Pock proximal algorithm [10, 1, 59] leveraging the closed-form expression of the proximity operator of the Kullback-Leibler divergence of Equation (18[1].
@@
Let $\boldsymbol{\mathsf{Y}}$ observations following the generalized time-varying autoregressive Poisson model described in Section 4.1. Then, the variational estimator $\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda)$ defined in Equation (41) satisfies the assumptions enunciated in Theorems 1 and 2. Hence, for any $N\in\mathbb{N}^{*}$, the data-driven robustified oracles $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}$ and $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}$ are asymptotically unbiased estimators of the estimation and prediction risks respectively.
@@
$\begin{split}\mathcal{E}^{\circ}(\boldsymbol{\mathsf{Y}};\lambda)=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda)-\overline{\boldsymbol{\mathsf{X}}}\rVert^{2}_{2},\quad\mathcal{P}^{\circ}(\boldsymbol{\mathsf{Y}};\lambda)=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda)\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})-\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\rVert^{2}_{2}\end{split}$ (42)
@@
Metric Minimized oracle $\log_{10}\alpha=2$ $\log_{10}\alpha=2.5$ $\log_{10}\alpha=3$ $\log_{10}\alpha=3.5$ $\log_{10}\alpha=4$ $\log_{10}\alpha=4.5$
$\mathsf{MMSE}\pm\mathsf{CI}$ Estimation error $\mathcal{E}^{\circ}$ $0.51\pm 0.02$ $0.57\pm 0.05$ $1.06\pm 0.20$ $1.23\pm 0.40$ $1.76\pm 0.52$ $1.37\pm 0.26$
Prediction error $\mathcal{P}^{\circ}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $1.21\pm 0.32$ $2.38\pm 2.00$ $1.89\pm 0.52$ $1.50\pm 0.27$
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $1.20\pm 0.33$ $1.40\pm 0.44$ $2.17\pm 0.68$ $1.53\pm 0.29$
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $2.03\pm 1.28$ $4.86\pm 4.10$ $7.42\pm 4.42$ $23.98\pm 17.82$
$\mathsf{Bias}$$\mathsf{Variance}$ Estimation error $\mathcal{E}^{\circ}$ 0.490.490.490.490.020.020.020.02 0.500.500.500.500.060.060.060.06 0.730.730.730.730.330.330.330.33 0.550.550.550.550.680.680.680.68 0.850.850.850.850.910.910.910.91 0.550.550.550.550.820.820.820.82
Prediction error $\mathcal{P}^{\circ}$ 0.490.490.490.490.040.040.040.04 0.500.500.500.500.090.090.090.09 0.660.660.660.660.550.550.550.55 0.820.820.820.821.561.561.561.56 0.770.770.770.771.111.111.111.11 0.480.480.480.481.021.021.021.02
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.500.500.030.030.030.03 0.530.530.530.530.070.070.070.07 0.700.700.700.700.500.500.500.50 0.560.560.560.560.840.840.840.84 0.910.910.910.911.261.261.261.26 0.570.570.570.570.960.960.960.96
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.500.500.030.030.030.03 0.510.510.510.510.080.080.080.08 0.720.720.720.721.311.311.311.31 0.950.950.950.953.913.913.913.91 0.990.990.990.996.436.436.436.43 1.471.471.471.4722.5122.5122.5122.51
@@ 4.3 Performance evaluation
For each scale parameter explored, performances are evaluated on a collection of $Q=20$ independent synthetic observations $\{\boldsymbol{\mathsf{Y}}^{(q)},\,q=1,\ldots,Q\}$ generated following the setup of Section 4.1. For $\mathcal{O}\in\{\mathcal{E}^{\circ},\mathcal{P}^{\circ},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\}$, the hyperparameters selected through the oracle strategy applied to the $q$th realization are denoted [^6] The dependency of $\lambda_{\mathcal{O}}$ in $\boldsymbol{\mathsf{Y}}$ is omitted for readability.
@@
Several performance criteria are considered. First, the Minimal Mean Squared Error over the $Q$ realizations:
@@
$\mathsf{Bias}=\left\lVert\langle\widehat{\boldsymbol{\mathsf{X}}}\rangle_{\mathcal{O}}-\overline{\boldsymbol{\mathsf{X}}}\right\rVert^{2}_{2},\quad\mathsf{Variance}=\frac{1}{Q}\sum_{q=1}^{Q}\left\lVert\widehat{\boldsymbol{\mathsf{X}}}\left(\boldsymbol{\mathsf{Y}}^{(q)};\lambda_{\mathcal{O}}^{(q)}\right)-\langle\widehat{\boldsymbol{\mathsf{X}}}\rangle_{\mathcal{O}}\right\rVert^{2}$ (46)
@@
Remark.
The number of realizations $Q$ over which the performance is computed has been chosen after observing that all aforementioned metrics together with their confidence intervals had stabilized after $Q=20$.
@@ 4.4 Results
Figure 2 compares the four oracle-based estimation strategies on one realization of generalized time-varying autoregressive Poisson synthetic observations for three values of the scale parameter, corresponding to low, medium and high noise levels. Synthetic observations and memory terms are displayed in the first row; the underlying ground truth reproduction coefficient and its estimates obtained using the four different oracles are plotted in blue on the second row; third (resp. fourth) row compares the prediction (resp. estimation) ground truth and data-dependent oracles, and the associated optimal hyperparameters.
@@ 5.1 Weekly scaled Poisson epidemiological model
The considered epidemiological model, briefly introduction in Section 1, in Equation (5), was initially proposed in [17] and states that, conditionally to past infection counts $\mathsf{Z}_{1},\ldots,\mathsf{Z}_{t-1}$, the number of new infections at time $t$, denoted $\mathsf{Z}_{t}$, follows a Poisson distribution of intensity equal to the product of the effective reproduction number at time $t$, $\mathsf{R}_{t}$, and the global infectiousness $\Phi_{t}(\boldsymbol{\mathsf{Z}})=\sum_{s\geq 1}\varphi_{s}\mathsf{Z}_{t-s}$ with $\varphi$ the serial interval distribution. [^8] The serial interval is the random delay between primary and secondary infections; its distribution encodes the typical time scales of the propagation. Although very accurate for a posteriori consolidated data [17], the daily version of this model appeared not suitable for real-time COVID-19 daily infection counts, an example of which is displayed in Figure LABEL:fig:daily, gray curve. Indeed, COVID-19 data are severely corrupted by administrative noise, taking the form of missing counts during week-ends and holidays, large cumulative counts on Mondays, pseudo-seasonalities, erroneous samples, to name but a few. This noise has two major effects: first, reported COVID-19 daily infections counts contain outlier values, notably during week-ends as illustrated in the gray curve in Figure LABEL:fig:daily; second, the variance of the reported counts reflects both the intrinsic variance of the propagation process and the additional variance induced by the fluctuating administrative delays and errors. To circumvent the presence of outlier samples, a classical strategy in epidemiology is to consider aggregated data [25, 13, 53], e.g., at the scale of the week as illustrated in Figure LABEL:fig:daily, black curve, which is far smoother and hence more realistic from an epidemiological point of view. To go further in the modeling of infection counts, the present work proposes to account for the increased variance observed in real COVID-19 infection counts through a scale parameter $\alpha>1$ constant in time. Up to the authors’ knowledge, scaled Poisson distributions have not yet been explored in the epidemiology literature, making this an original contribution. Altogether, the proposed weekly scaled Poisson epidemiological model writes
@@
where time instants $t$ corresponds to weeks and the serial interval distribution used to compute $\Phi_{t}(\boldsymbol{\mathsf{Z}})$ is weekly discretized. The weekly discretized COVID-19 serial interval distribution can be obtained by coarsening the daily discretized distribution provided in [9, 66] into a weekly distribution, e.g., by performing an integration over one-week windows using the left rectangle method with a one-day integration step. Under Model (47), the expected number of infections at week $t$ is $\Phi_{t}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{t}$, unchanged compared to the standard Poisson model of Equation (5), but the variance is $\alpha\Phi_{t}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{t}$. If $\alpha>1$, then the variance is larger by a factor $\alpha$ than the standard Poisson model variance.
@@ 5.2 Data-driven reproduction number estimation strategy
leveraging the robustified unbiased prediction risk estimate of Equation (35), whose accuracy and robustness have been demonstrated on synthetic data in Section 4.
@@ 5.3 Experimental setup
(a) France
(b) India
(c) Canada
(d) Argentina
Figure 4: Data-driven estimation of the reproduction number of COVID-19 from weekly infection counts. Top rows: weekly aggregated infection counts. Middle rows: estimated effective reproduction number with data-driven hyperparameter selection based on the robustified $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$. Bottom rows: robustified autoregressive unbiased prediction risk estimate averaged over $N=10$ realizations of the Monte Carlo vector with $95\%$ Gaussian confidence interval for a logarithmically spaced range of regularization parameters and minimizing regularization parameter obtained through grid search.
@@ 6 Conclusion and perspectives
A time-varying autoregressive model generalizing state-of-the-art viral epidemics models has been considered. The estimation of the parameters of the model, namely the time-varying reproduction coefficient, has been framed as a nonstandard, highly nonstationary, inverse problem. The proposed rigorous mathematical formulation enabled to leverage the efficient and versatile variational framework to design accurate reproduction coefficient estimators which are robust to high noise levels in the observations. The major contribution consisted in the design of asymptotically unbiased risk estimates, suited to the generalized autoregressive Poisson model, which are then plugged into an oracle strategy for fully data-driven fine-tuning of the hyperparameters of the variational estimator, removing the major obstacle to its practical use. The resulting data-driven time-varying estimation strategy has been assessed through intensive Monte Carlo simulations on synthetic data. Taking advantage of the proposed extended mathematical framework, the standard epidemiological model for viral epidemics has been enriched to account for the low quality of COVID-19 data, leading to a novel weekly scaled Poisson model. Finally, the data-driven time-varying estimation procedure is shown to yield very consistent estimation of the COVID-19 reproduction number in various countries and pandemic stages despite the low quality of reported data, demonstrating its practical applicability for epidemic monitoring in a crisis context. The data-driven nature of the proposed epidemiological indicator estimation strategy constitutes a major asset for its dissemination beyond signal processing as its use requires no expert knowledge.
Future work will consist in, first, extending the unbiased risk estimates to other noise models, e.g., to additive Gaussian and multiplicative Gamma noises inspiring from the generalized Stein estimators [23], in order to further enrich the proposed framework and to be able to extend the developed methodology to other models, in epidemiology or beyond. Second, as in epidemiology the memory functions are often parametric, with parameters encapsulating the pathogen transmission characteristics, the proposed framework will be leveraged to perform simultaneously the extraction of the memory functions parameters and the fine-tuning of the estimator hyperparameters.
@@
7 Acknowledgements
B. Pascal is supported by ANR-23-CE48-0009 “OptiMoCSI” and S. Vaiter by ANR-18-CE40-0005 “GraVa”. The authors thank Patrice Abry for fruitful discussions.